A baseball bat with 100. J of energy strikes a ball with a mass of 150. g. The energy transfer is only 20.0 % efficient. Calculate the initial speed of the ball.


1
Expert's answer
2020-10-29T07:01:22-0400

"A = Q+E_k"

"E_k = \frac{20*A}{100}=20\text{ J}"

"E_k=\frac{mV^2}{2};\ V=\sqrt{\frac{E_k*2}{m}}=\sqrt{\frac{20*2}{0.150}}= \approx16.33\ \frac{m}{s}"

Answer:"V= \approx16.33\ \frac{m}{s}"